OpenStudy (anonymous):

write an equation of a line that passes that passes through the point (7, 3) and is parallel to the line y = -2/3x + 3

jimthompson5910 (jim_thompson5910):

y = -2/3x + 3 has a slope of -2/3 since it's in the form y = mx+b (m is your slope)

jimthompson5910 (jim_thompson5910):

any line parallel to y = -2/3x + 3 will also have a slope of -2/3 parallel lines have equal slopes

jimthompson5910 (jim_thompson5910):

so you know the equation for the answer looks like y = (-2/3)x + b where b is some fixed number. At this point, b is unknown, but you can find the value of b. You plug in (x,y) = (7,3) which is your given point and solve for b. I'll let you do this part.